# Selective labeling with false discovery rate control

### v1

@@ 3.1 Conformal Labeling
Suppose calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are i.i.d. Let $\alpha\in(0,1)$ be the target FDR level, and suppose the selection set $\mathcal{R}$ is determined by Algorithm 1 applied at the target FDR level $\alpha$. Define $p=\mathbb{E}[H_{j}^{0}]$, the probability that a test sample $(X_{n+j},Y_{n+j})$ is incorrectly predicted. Then, the FDR of the selection set $\mathcal{R}$ satisfies:
$\text{FDR}\leq\left[1-(1-p)^{n+1}\right]\alpha\leq\alpha.$
3.2 Choice of uncertainty score
As shown in prior work (Jin & Candès, 2023; Gui et al., 2024; Bai et al., 2025b), the statistical power of this method depends on the quality of the uncertainty score. In particular, a score that better separates correct from incorrect predictions directly increases statistical power (see Proposition 7 of Jin & Candès (2023)). To deliver practical recommendations, we empirically compare several uncertainty scores by visualizing their resulting $p$-value distributions and measuring the final statistical power employed with our method. We utilize a pre-trained ResNet-34 model on the ImageNet dataset, with three uncertainty scores: maximum softmax probability (MSP) (Hendrycks & Gimpel, 2016), energy score (Liu et al., 2020), and DOCTOR-$\alpha$ score ($D_{\alpha}$) (Granese et al., 2021). We give an overview of these score functions in Appendix C. The results in Figure 1 show that both MSP and $D_{\alpha}$ score provide a clear distinction between correct and incorrect predictions, thus achieving high statistical power. However, the energy score fails to provide this separation, leading to low power. Given the comparable power performance of MSP and $D_{\alpha}$ score, we will use the more computationally efficient MSP in our main experiments.
@@ 4 Experiments
In this section, we evaluate the effectiveness of Conformal Labeling on image labeling, text labeling, and LLM QA tasks with various models. We find that it achieves tight FDR control and high power, indicating that AI models can label a large proportion of data with high quality. We also conduct comprehensive ablation studies to provide practical guidance for applying our method.
@@ 4.1 Experimental setup
Tasks and datasets.
We evaluate the effectiveness of Conformal Labeling across three labeling tasks, including image labeling, text labeling, and LLM question answering (QA). In appendix G, we also demonstrate how to extend Conformal Labeling to regression tasks. For the LLM QA task, the goal is to identify subsets of questions that LLMs can answer correctly. We employ common benchmark datasets for evaluations in each labeling task. For image classification, we use ImageNet (Deng et al., 2009) and its variant ImageNet-V2 datasets (Recht et al., 2019). For text labeling, we adopt two benchmark datasets. The first is Stance on Global Warming (Luo et al., 2021), which provides annotations ($Y_{i}\in\{\text{agree},\text{neutral},\text{disagree}\}$) to judge whether a headline agrees that global warming is a serious concern. The second is Misinformation (Gabriel et al., 2022), which contains binary annotations ($Y_{i}\in\{\text{misinfo},\text{real}\}$) for identifying whether a given text contains misinformation. For the LLM QA task, we evaluate our method on MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024b) datasets.
Table 1: Performance of Conformal Labeling on three labeling tasks. $\uparrow$ indicates that a larger value is better. We evaluate on Image Labeling (ImageNet, ImageNet-V2), text Labeling (Stance on global warming, Misinformation), and LLM QA (MedMCQA, MMLU) tasks. We report results for Conformal Labeling at $\alpha=0.1$ and compare Conformal Labeling against two baselines: (i) a naive approach of labeling instances whose uncertainty score $\mathcal{S}_{n+j}\leq 0.1$ with AI models, and (ii) labeling the entire dataset with AI models. The results show that Conformal Labeling consistently achieves tighter FDR control across all datasets and models compared with the baseline.
Task Dataset Model Conformal Labeling ($\alpha$= 0.1) Naive ($\mathcal{S}\leq 0.1$) AI only
FDR % Power %($\uparrow$) Ratio %($\uparrow$) FDR % Power %($\uparrow$) Ratio %($\uparrow$) Error %($\downarrow$)
Image ImageNet ResNet-34 9.97 80.01 58.67 4.79 63.09 43.71 26.71
DenseNet-161 9.99 85.03 65.56 5.58 72.08 52.98 22.89
ResNeXt50 10.00 86.06 66.83 6.08 75.57 56.21 22.39
CLIP-VIT-B/32 9.98 46.04 27.47 5.53 28.65 16.28 40.35
ImageNet-V2 ResNet-34 10.00 61.99 37.87 7.25 56.03 33.15 39.03
DenseNet-161 9.83 66.67 43.39 9.39 65.63 42.45 34.88
ResNeXt50 9.95 67.86 44.75 10.17 68.71 45.38 34.08
CLIP-VIT-B/32 9.96 34.56 18.10 7.86 25.83 13.18 47.78
Text Stance Llama-3.1-8B-Instruct 9.77 19.02 9.42 24.43 54.94 31.39 52.04
Qwen3-32B 9.56 10.70 7.25 14.42 16.64 11.11 36.52
Qwen2.5-72B-Instruct 9.71 26.97 17.84 26.27 74.49 59.01 35.09
Misinformation Llama-3.1-8B-Instruct 9.88 7.31 5.81 18.38 66.17 55.25 24.28
Qwen3-32B 9.91 49.21 37.48 10.77 52.27 40.03 24.08
Qwen2.5-72B-Instruct 9.58 44.36 34.81 17.81 87.13 74.72 21.69
QA MedMCQA Llama-3.1-8B-Instruct 9.70 31.44 18.90 15.01 46.74 29.53 40.35
Qwen3-32B 9.75 49.80 33.27 13.79 65.27 45.36 33.44
Llama-3.1-70B-Instruct 9.95 69.67 49.59 4.52 48.92 32.79 28.90
MMLU Llama-3.1-8B-Instruct 9.99 58.25 37.47 8.47 53.72 33.96 35.72
Qwen3-32B 10.00 82.96 65.22 8.13 78.47 60.40 21.43
Llama-3.1-70B-Instruct 9.96 88.20 72.10 4.18 67.94 52.17 18.24
@@
We conduct extensive experiments on various open-sourced AI models. For image classification, we utilize three well-established deep image classifiers: ResNet-34 (He et al., 2016), DenseNet-161 (Huang et al., 2018), and ResNext50 (Xie et al., 2017). Additionally, we employ the Vision-Language Model CLIP (Radford et al., 2021), which is based on a Vision Transformer architecture (ViT-B/32) (Dosovitskiy et al., 2021). The above classifiers are provided by TorchVision (Paszke et al., 2019). For text labeling, we employ three LLMs: Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Qwen3-32B, Qwen2.5-72B-Instruct (Qwen et al., 2025). For LLM QA tasks, we employ five LLMs: Qwen3-8B (Yang et al., 2025), Qwen3-14B, Qwen3-32B, Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct. The above LLMs are provided by Hugging Face.
@@
We evaluate Conformal Labeling with two baseline methods: using AI models to label test instances with uncertainty scores $\mathcal{S}_{n+j}\leq 0.1$, and applying AI predictions to the entire test dataset. We compare Conformal Labeling’s selection procedure against BH (Benjamini & Hochberg, 1995), Storey-BH (Storey, 2002), and Quantile-BH procedures (Benjamini et al., 2006). We evaluate the performance of our method and baseline using the following metrics: (1) FDR, the expected proportion of incorrect labels in the selected set $\mathcal{R}$; (2) Power, the proportion of correctly labeled instances selected; (3) AI-labeled ratio, the number of data labeled by the AI models divided by the combined size of the calibration and test datasets.
@@
To ensure the reliability of our results, we repeat each experiment 1000 times and report the average result. We randomly select 10% of the data as the calibration dataset. For all the experiments, we use the maximum softmax probability (MSP) as the uncertainty score function. In the LLM QA task, we adopt the standard multiple-choice evaluation pipeline: given a question and candidate answers (e.g., A, B, C, or D), the model estimates the probability of each option by extracting the logits corresponding to the option tokens and applying a softmax transformation, with the predicted label taken as the option with the highest probability. For text labeling, we reformulate each sample into a multiple-choice format (e.g., “positive,” “negative,” or “neutral”), enabling the same probability-extraction procedure as in the LLM QA task. More details of implementation are provided in Appendix F. We provide the code for reproducing our main experiments in this anonymous repository.
4.2 Experiments results
@@
In table 1, we present the performance of Conformal Labeling against two baselines on three different labeling tasks: image labeling, text labeling, and LLM QA. A salient observation is that across all the labeling tasks and all the model architectures, Conformal Labeling successfully controls the FDR at or below the target FDR level. In comparison, both baseline methods lack FDR control, resulting in substantial labeling errors that compromise label quality when AI models are inaccurate. It is worth noting that the FDR is tightly controlled: for $\alpha=10\%$, most experiments yield FDRs below $9.9\%$, with the largest deviation at $9.56\%$. This tight FDR control directly leads to high selection power. For example, on MMLU with Qwen3-32B at $\alpha=10\%$, Conformal Labeling achieves a power of $82.96\%$, labeling $65.22\%$ of the dataset with the AI model. Overall, empirical results show that Conformal Labeling consistently achieves tight FDR control with high power across different datasets and models.
Higher prediction accuracy enables better selection results.
The performance of Conformal Labeling depends heavily on the underlying prediction accuracy, which is influenced by model capacity and dataset difficulty. We evaluate Conformal Labeling with Qwen3-8B, Qwen3-14B, and Qwen3-32B, whose increasing scales provide greater capacity. MMLU-Pro, more challenging than MMLU, results in lower prediction accuracy. Figure 2 presents the evaluation results. In all cases, the realized FDR remains below and close to $\alpha=0.1$. Higher accuracy—achieved with stronger models or easier datasets—boosts power and the AI-labeled ratio. Specifically, for any given model, performance is superior on MMLU compared to MMLU-Pro. Similarly, for each dataset, larger models yield greater power and AI-labeled ratios. Overall, higher prediction accuracy leads to better selection results by achieving higher power and AI-labeled ratio with controlled FDR.
@@
The size of the calibration set plays a crucial role in constructing reliable conformal $p$-values. To study the effect of $|D_{\mathrm{cal}}|$ on FDR and power, in Figure 3, we label {5%, 10%, 20%} of the unlabeled dataset as the calibration dataset. Our results demonstrate that Conformal Labeling is robust to calibration set size: even with a 5% calibration ratio, the FDR remains controlled with low standard deviation. Increasing the calibration ratio reduces the variance of both FDR and power, although the improvement from 10% to 20% is negligible. Based on this trade-off between variance reduction and labeling cost, we use a 10% calibration ratio for all the experiments. In summary, while Conformal Labeling is robust to calibration set sizes, larger calibration sets reduce the variance of FDR and power, thereby enhancing selection stability.
Figure 3: Performance of Conformal Labeling with varying calibration set sizes on the MedMCQA dataset. The top row corresponds to the results from Qwen3-8B and the bottom row to those from Qwen3-32B; each column corresponds to a value of calibration ratio. Shaded regions indicate one standard deviation around the average. The results show that a large calibration set slightly reduces the variance of FDR and power, leading to a more robust selection outcome. Overall, our method is robust to changes in the calibration set size.
Figure 4: Albation study on the selection procedure. The ablation study is conducted on the MMLU dataset with Qwen3-32B. We substitute Conformal Labeling’s selection procedure with alternative selection procedures, including BH, Storey-BH, and Quantile-BH procedures. The figure shows the Conformal Labeling’s selection procedure consistently achieves the tightest FDR control and the highest power among all the selection procedures.
Ablation study on the selection procedure.
To validate the effectiveness of Conformal Labeling’s selection procedure, we conduct an ablation study substituting it with alternative selection procedures, including BH, Storey-BH (Storey, 2002), and Quantile-BH (Benjamini et al., 2006) procedures (see Appendix D for details). The hyperparameters for the Storey-BH and Quantile-BH procedures are chosen using a bootstrap method detailed in the Appendix E. An ideal selection procedure aims to maximize power while keeping the FDR below the desired level.
The results of this ablation study are presented in Figure 4. While all the selection procedures successfully control the FDR below the desired level, Conformal Labeling’s selection procedure (red line) provides the tightest control, with realized FDR consistently closest to the desired level (dashed blue line). In comparison, alternative selection procedures are more conservative, especially at low target FDR levels, which reduces power. Overall, this ablation study highlights the superiority of Conformal Labeling’s selection procedure.

### v2

@@ 3.1 Conformal Labeling
Suppose the calibration samples $(X_{i},Y_{i})_{i=1}^{n}$ and test samples $(X_{n+j},Y_{n+j})_{j=1}^{m}$ are i.i.d. Let $\alpha\in(0,1)$ be the target FDR level, and suppose the selection set $\mathcal{R}$ is determined by Algorithm 1 applied at the target FDR level $\alpha$. Define $p=\mathbb{E}[H_{j}^{0}]$, the probability that a test sample $(X_{n+j},Y_{n+j})$ is incorrectly predicted. Taking expectation over the randomness of the calibration and test data, the FDR of the selection set $\mathcal{R}$ satisfies:
$\mathrm{FDR}\leq\left[1-(1-p)^{n+1}\right]\alpha\leq\alpha.$
@@
3.2 Choice of uncertainty score
As shown in prior work (Jin & Candès, 2023; Gui et al., 2024; Bai et al., 2025), the statistical power of this method depends on the quality of the uncertainty score. In particular, a score that better separates correct from incorrect predictions directly increases statistical power (see Proposition 7 of Jin & Candès (2023)). To deliver practical recommendations, we empirically compare several uncertainty scores by visualizing their resulting $p$-value distributions and measuring the final statistical power employed with our method in Figure 1. We utilize a pre-trained ResNet-34 model on the ImageNet dataset, with three uncertainty scores: maximum softmax probability (MSP) (Hendrycks & Gimpel, 2016), energy score (Liu et al., 2020), and DOCTOR-$\alpha$ score ($D_{\alpha}$) (Granese et al., 2021). We give an overview of these score functions in Appendix B. The results show that both MSP and $D_{\alpha}$ score provide a clear distinction between correct and incorrect predictions, thus achieving high statistical power. However, the energy score fails to provide this separation, leading to low power. Given the comparable power performance of MSP and $D_{\alpha}$ score, we will use the more computationally efficient MSP in our main experiments. In Appendix I.2, we provide a detailed comparison of the power and FDR achieved by these score functions.
@@ 4 Experiments
In this section, we evaluate the effectiveness of Conformal Labeling on image labeling, LLM QA, and LLM open-ended generation tasks with various models. We find that it achieves tight FDR control and high power, indicating that AI models can label a large proportion of data with bounded error. In addition, we conduct additional analyses to provide practical guidance for applying our method.
@@ 4.1 Experimental setup
Datasets.
We conduct experiments on image labeling, LLM QA, and LLM open-ended generation tasks. For image labeling, we use ImageNet (Deng et al., 2009) and ImageNet-V2 (Recht et al., 2019). For LLM QA task, we use MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024c). For LLM open-ended generation task, we use MMLU-Redux (Gema et al., 2025), MATH-500 (Lightman et al., 2023), MATH-L5 (Hendrycks et al., 2021), Zebra-Logic (Lin et al., 2025), and $\text{HumanEval}^{+}$ (Liu et al., 2023) to evaluate Conformal Labeling’s performance in general, math reasoning, text reasoning and coding tasks respectively. We also show how to extend our method to regression tasks in Appendix H.
Table 1: Annotation performance of Conformal Labeling on image Labeling and QA tasks. We compare Conformal Labeling against three baselines: (1) selection with guaranteed risk control (SGR), (2) FDR search heuristic method, and (3) labeling the entire dataset with AI models. We set $\alpha=0.1$ for Conformal Labeling and FDR search, and $\alpha=0.1$, $\epsilon=0.2$ for SGR. Bold numbers are superior results.
Task Dataset Model Conformal Labeling SGR FDR search AI only
FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) Error (%)
Image ImageNet ResNet-34 9.97 79.89 65.04 8.91 77.31 62.21 9.85 79.63 64.75 26.71
DenseNet-161 9.92 84.89 72.68 8.92 82.68 70.00 9.83 84.71 72.45 22.89
CLIP-VIT-B/32 9.84 45.63 30.21 8.17 39.66 25.79 9.77 45.41 30.04 40.35
ImageNet-V2 ResNet-34 9.92 61.78 41.85 6.69 52.20 34.19 9.78 61.51 41.62 39.04
DenseNet-161 9.98 67.07 48.57 6.82 54.18 37.98 9.84 66.67 48.21 34.88
CLIP-VIT-B/32 9.99 35.52 20.73 8.17 6.59 3.67 9.51 33.61 19.55 47.78
QA MedMCQA Llama-3.1-8B-Instruct 9.80 31.72 21.20 8.54 2.03 1.27 9.39 30.41 20.27 40.35
Qwen3-32B 9.81 49.85 37.05 8.08 2.94 2.09 9.55 48.60 36.05 33.44
Llama-3.1-70B-Instruct 9.86 69.37 54.83 6.43 44.34 33.67 10.58 71.15 56.67 28.90
MMLU Llama-3.1-8B-Instruct 9.85 57.13 40.64 7.00 48.00 33.13 58.94 76.77 68.74 35.91
Qwen3-32B 10.00 82.95 72.44 8.18 78.14 66.88 9.84 82.56 71.98 21.43
Llama-3.1-70B-Instruct 9.97 88.25 80.17 8.21 84.00 74.85 9.90 88.10 79.98 18.24
Table 2: Annotation performance of Conformal Labeling on open-ended generation tasks. We use the logits-based uncertainty score. We set $\alpha=0.05$ for Conformal Labeling and FDR search, and $\alpha=0.05$, $\epsilon=0.2$ for SGR. Bold numbers indicate superior results.
Dataset Model Conformal Labeling SGR FDR search AI only
FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) FDR (%) Power (%) Ratio (%) Error (%)
MMLU-Redux Qwen3-4B-Instruct-2507 4.84 33.62 28.46 43.17 9.06 7.71 4.22 29.87 25.11 20.17
Qwen3-4B-Thinking-2507 4.95 24.64 21.46 42.56 4.08 3.61 5.85 23.77 20.62 20.79
DeepSeek-R1-Distill-Qwen-32B 4.92 29.19 26.34 51.35 6.89 6.30 5.80 27.48 24.70 15.07
MATH500 Qwen3-4B-Instruct-2507 4.93 86.73 83.88 3.18 52.83 50.72 4.88 84.94 82.10 8.20
Qwen3-4B-Thinking-2507 4.07 99.80 99.78 2.50 57.25 57.11 4.01 99.69 99.67 4.00
DeepSeek-R1-Distill-Qwen-32B 4.18 8.28 8.41 38.70 0.31 0.61 29.83 10.32 10.63 7.00
MATH-L5 Qwen3-4B-Instruct-2507 4.84 69.54 66.90 0.55 2.90 2.85 5.14 68.89 66.25 8.88
Qwen3-4B-Thinking-2507 4.64 48.33 48.45 0.35 1.96 1.95 4.72 57.85 57.94 5.27
DeepSeek-R1-Distill-Qwen-32B 4.69 65.31 64.69 1.42 13.86 13.71 6.10 68.66 68.04 6.38
Zebra-Logic Qwen3-4B-Instruct-2507 4.76 47.80 39.94 0.75 6.52 5.46 4.23 42.36 35.18 20.90
Qwen3-4B-Thinking-2507 4.95 67.83 63.73 2.45 32.11 30.02 3.05 50.24 46.28 11.00
DeepSeek-R1-Distill-Qwen-32B 4.78 0.01 0.02 69.33 0.03 0.03 67.70 0.05 0.32 31.40
HumanEval+ Qwen3-4B-Instruct-2507 4.76 72.46 71.06 0.44 5.60 5.38 5.09 81.86 80.08 7.32
Qwen3-4B-Thinking-2507 2.44 100.00 100.00 1.61 14.73 14.79 2.44 100.00 100.00 2.44
DeepSeek-R1-Distill-Qwen-32B 0.61 100.00 100.00 0.64 39.48 39.49 0.62 100.00 100.00 0.61
@@
We conduct extensive experiments on various open-sourced AI models. For image labeling, we use ResNet-34 (He et al., 2016), DenseNet-161 (Huang et al., 2018), and a Vision-Language Model CLIP (Radford et al., 2021), which is based on a pre-trained ViT-B/32 (Dosovitskiy et al., 2021). For LLM QA tasks, we employ Llama-3.1-8B-Instruct (Dubey et al., 2024), Qwen3-32B (Yang et al., 2025), and Llama-3.1-70B-Instruct. For LLM open-ended generation tasks, we use Qwen3-4B-Instruct-2507, Qwen3-4B-Thinking-2507, and DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025). The specific sampling temperature and other hyperparameters for LLMs are detailed in Appendix F. While the theoretical guarantee of Conformal Labeling is model-agnostic, we show in Appendix I.3 that higher AI model accuracy leads to a larger selection size.
Uncertainty Scores.
For the image labeling task, we use the MSP. For the LLM QA task, we also adopt MSP, where the estimated probability of each option is obtained by applying a softmax over the logits corresponding to the option tokens. For the LLM open-ended generation task, we adopt two complementary score functions that serve as analogues of MSP: a logits-based score for the white-box case and a verbalized score for the black-box case. We provide details for these score functions in Appendix E.
@@
We evaluate Conformal Labeling against four baseline methods: (1) an FDR-search heuristic baseline; (2) selection with guaranteed risk control (SGR) (Geifman & El-Yaniv, 2017); (3) labeling test instances with AI models when the uncertainty score satisfies $\mathcal{S}_{n+j}\leq\alpha$; and (4) applying AI predictions to the entire test dataset. We provide a detailed introduction to these baseline methods in Appendix G. We compare Conformal Labeling’s selection procedure against BH (Benjamini & Hochberg, 1995), Storey-BH (Storey, 2002), and Quantile-BH procedures (Benjamini et al., 2006). We evaluate the performance of selective labeling using the following metrics: (1) FDR; (2) Power; (3) AI-labeled ratio, the proportion of data in the test dataset that is labeled by AI models.
@@
Each experiment is repeated 1000 times, and we report the mean FDR, power, and AI-labeled ratio. More details of implementation are provided in Appendix F. We provide the code for reproducing our main experiments in this anonymous repository.
4.2 Results
@@
Table 1 reports the annotation performance of Conformal Labeling on image labeling and LLM QA tasks. Across all the benchmarks and model architectures, Conformal Labeling tightly controls the FDR while labeling a substantial proportion of data with AI models. For example, using Llama-3.1-70B-Instruct on MedMCQA, Conformal Labeling labels 54.83% of test instances with the AI model while controlling the FDR at 9.86%. In comparison, the FDR search baseline fails to control FDR, while SGR only labels 33.67% of test instances. Conformal Labeling improves the AI-labeled ratio by over 20% compared to SGR, while controlling the FDR. The tightest FDR control and highest power are also consistently observed in other datasets and models. Overall, Conformal Labeling achieves the best cost-accuracy tradeoff.
Conformal Labeling outperforms baselines in the open-ended generation task.
Table 2 reports the annotation performance of Conformal Labeling on open-ended generation task. The results show that Conformal Labeling reliably controls the FDR on open-ended generation tasks, whereas baselines may incur substantially violated FDR. For example, on MMLU-Redux with Qwen3-4B-Thinking-2507, Conformal Labeling controls the FDR at 4.95% while labeling 21.46% of instances, whereas SGR incurs an FDR of 42.56% and labels only 3.61% of data, and FDR search also exceeds the target FDR with an FDR of 5.85%. Besides, when baselines control the FDR below the target level, Conformal Labeling consistently achieves higher power than the baselines. Overall, Conformal Labeling achieves the best cost-accuracy tradeoff among all the compared methods.
(a) MMLU-Redux
(b) MATH-L5
(c) Zebra-Logic
Figure 2: Power of Conformal Labeling with differnet score functions. The experiments are conducted with Qwen3-4B-Instruct-2507 on three benchmarks. Shaded areas indicate standard deviations. We present the corresponding realized FDR in Figure 5.
Logits-based uncertainty score outperforms verbalized uncertainty score in open-ended generation task.
Figure 2 compares logits-based and verbalized score functions under varying target level $\alpha$. Under different score functions, Conformal Labeling consistently controls the FDR below the target level, while using logits-based score function results in higher power. For example, at $\alpha=0.07$, the logits-based score achieves over 50% power, whereas the verbalized score attains power of approximately 25%, corresponding to an absolute power improvement of over 25%. We also plot the reliability diagrams of these score functions in Figure 6, which shows that logits-based scores achieve lower expected calibration error (ECE) (Naeini et al., 2015) in most datasets, indicating superior reliability. We also plot the $p$-value distribution under these two score functions in Figure 7, which demonstrates that the logits-based uncertainty score provides a clearer separation between incorrect and correct labels. Together, these results demonstrate that the logits-based score provides more reliable uncertainty estimates than the verbalized score, leading to higher power.
Conformal Labeling outperforms existing selection procedures.
We conduct an ablation study on the selection procedure by fixing $p$-values and varying only the selection procedure. We compare Conformal Labeling’s selection procedure with three baselines: Storey-BH (Storey, 2002), Quantile-BH (Benjamini et al., 2006), and the standard BH procedure. Unlike Conformal Labeling that estimates $\pi_{0}$ from the calibration set, Storey-BH and Quantile-BH procedures estimate $\pi_{0}$ from the $p$-values. We provide a detailed introduction to these procedures in Appendix C. Besides, while Conformal Labeling’s selection procedure requires no hyperparameter tuning, Storey-BH and Quantile-BH involve hyperparameters that must be carefully specified. In this work, we set hyperparameters for these procedures using a bootstrap method (Storey, 2002) detailed in Appendix D.
The comparison results in Table 3 show that our method can achieve much higher power than other selection procedures, with more precise FDR control. For instance, on MATH-L5 with Qwen3-4B-Instruct-2507, Conformal Labeling achieves a power of 94.25% at $\alpha=0.1$. In comparison, the BH, Storey-BH, and Quantile-BH procedures yield power of 0.09%, 61.42%, and 29.00%, respectively. Conformal Labeling’s selection procedure improves power by over 30% compared to these baselines. The highest power is also consistently observed in other datasets and models. In summary, Conformal Labeling’s selection procedure performs the best among all the compared selection procedures.
Table 3: Comparison of different selection procedures on three labeling tasks. We compare Conformal Labeling’s selection procedure against three baselines: (1) BH procedure, (2) Storey-BH procedure, and (3) Quantile-BH procedure. We report results at $\alpha=0.1$ for all four selection procedures. Bold numbers are superior results.
Task Dataset Model CL BH Storey-BH Quantile BH
FDR % Power % FDR % Power (%) FDR (%) Power (%) FDR (%) Power (%)
Image ImageNet ResNet-34 9.97 80.01 2.63 46.85 8.64 74.57 7.57 71.79
DenseNet-161 9.99 85.03 2.27 44.55 8.34 81.17 7.33 78.40
CLIP-VIT-B/32 9.98 46.04 4.08 21.70 8.52 41.02 8.05 39.23
ImageNet-V2 ResNet-34 10.00 61.99 3.84 37.60 9.65 59.84 8.53 57.35
DenseNet-161 9.83 66.67 3.39 28.39 9.62 65.45 8.44 61.66
CLIP-VIT-B/32 9.96 34.56 4.85 14.43 9.20 32.25 8.35 28.52
QA MedMCQA Llama-3.1-8B-Instruct 9.70 31.44 3.68 12.89 7.43 25.37 6.33 22.11
Qwen3-32B 9.75 49.80 2.71 10.99 6.99 34.89 6.34 31.27
Llama-3.1-70B-Instruct 9.95 69.67 2.79 29.67 8.48 64.09 6.74 58.06
MMLU Llama-3.1-8B-Instruct 9.99 58.25 3.40 29.49 7.31 49.74 7.17 49.16
Qwen3-32B 10.00 82.96 1.37 10.51 7.33 74.57 6.25 70.37
Llama-3.1-70B-Instruct 9.96 88.20 1.61 18.74 6.95 80.08 5.96 76.77
OE MMLU-Redux Qwen3-4B-Instruct-2507 9.96 67.62 1.85 5.52 7.50 50.52 5.15 36.72
Qwen3-4B-Thinking-2507 9.88 66.37 1.59 1.71 4.93 26.86 3.77 15.58
DeepSeek-R1-Distill-Qwen-32B 9.84 77.79 1.40 1.80 4.50 29.21 3.44 19.43
MATH-L5 Qwen3-4B-Instruct-2507 7.97 94.25 0.99 0.09 5.38 61.42 3.13 29.00
Qwen3-4B-Thinking-2507 5.16 95.69 0.68 0.02 1.69 5.66 0.91 1.61
DeepSeek-R1-Distill-Qwen-32B 6.09 95.28 0.78 0.04 3.67 41.55 1.93 9.64
Zebra-Logic Qwen3-4B-Instruct-2507 10.00 72.70 1.87 4.68 9.45 71.15 6.40 50.41
Qwen3-4B-Thinking-2507 9.17 93.46 1.16 0.50 6.98 71.90 3.79 45.85
DeepSeek-R1-Distill-Qwen-32B 8.65 0.14 3.69 0.01 2.80 0.11 3.02 0.01
@@
To study the effect of $|D_{\mathrm{cal}}|$ on FDR and power, in Figure 3, we label {5%, 10%, 20%} of the unlabeled dataset as the calibration set. Our results show that Conformal Labeling is robust to calibration set size: even with a 5% calibration ratio, the FDR remains controlled with low standard deviation. Increasing the calibration ratio reduces the variance of both FDR and power, although the improvement from 10% to 20% is negligible. Based on this trade-off between variance reduction and labeling cost, we use a 10% calibration ratio for image labeling and LLM QA tasks. Overall, while Conformal Labeling is robust to calibration set sizes, larger calibration sets reduce the variance of FDR and power.
Figure 3: Performance of Conformal Labeling with varying calibration set sizes on MedMCQA with Qwen3-32B. The shaded areas represent standard deviations.
Due to space constraints, we defer additional analyses to the appendix. In particular, Appendix I.5 compares our method against FDR search and selective prediction based on calibrated confidence baselines; Appendix I.6 examines whether the selected labels can be used to help train a model with improved accuracy on a synthetic dataset; and Appendix I.7 evaluates Conformal Labeling’s sensitivity to prompt designs in LLM QA tasks.